OpenStudy (anonymous):

need a solutions please. thanks in adavnce x^2-3x-1=0

OpenStudy (callisto):

x= 3.30 or x=-0.303, cor to 3 s.f

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

Complete The Square

OpenStudy (anonymous):

(3+/-sqrt(9-4(1)(-1))/2 (3+/- sqrt(13) )/2

OpenStudy (callisto):

calculator!

OpenStudy (anonymous):

\[\left\{x\to \frac{1}{2} \left(3-\sqrt{13}\right),x\to \frac{1}{2} \left(3+\sqrt{13}\right)\right\} \]

OpenStudy (anonymous):

(x-1.5)^2=3.25

OpenStudy (anonymous):

thanks all